# thanks: These two authors contribute equally in this workthanks: These two authors contribute equally in this work

### v1

@@
Abstract
@@
Since the very recent discovery of unconventional superconductivity in twisted WSe2 homobilayers at filling $\nu=-1$, considerable interests arise in revealing its mechanism. In this paper, we developed a three-band tight-binding model with non-trivial band topology by direct Wannierization of the low-energy continuum model. Incorporating both onsite Hubbard repulsion and next-nearest-neighbor attraction, we then performed a mean-field analysis of the microscopic model and obtained a phase diagram qualitatively consistent with the experiment results. For zero or weak displacement field, the ground state is a Chern number $C=\pm 2$ topological superconductor in the Altland-Zirnbauer A-class (breaking time-reversal but preserving total $S_{z}$ symmetry) with inter-valley pairing dominant in $d_{xy}\mp id_{x^{2}-y^{2}}$-wave (mixing with a subdominant $p_{x}\pm ip_{y}$-wave) component. For a relatively strong displacement field, the ground state is a correlated insulator with $120^{\circ}$ antiferromagnetic order. Our results provide new insights into the nature of the twisted WSe2 systems and suggest the need for further theoretical and experimental explorations.
@@
Introduction.—In recent years, twisted van der Waals moiré superlattices garnered significant interests following the groundbreaking discovery of unconventional superconductivity (SC) [1, 2, 3] in magic-angle twisted bilayer graphene [4]. Beyond graphene-based systems, twisted transition metal dichalcogenides (TMDs) bilayers [5, 6] are regarded as a promising alternative platform for investigating many-body physics. Due to their high tunability, twisted TMDs can host a wide variety of exotic phases, including correlated insulators [7, 8, 9, 10], integer and fractional quantum anomalous Hall states [11, 12, 13, 14], and also integer and fractional quantum spin Hall states [15]. However, the experimental realization of SC in twisted TMD systems has remained elusive [9], though many theoretical works have suggested that SC should develops upon doping the correlated insulators [16, 17, 18, 19, 20, 21, 22, 23]. Very recently, two independent studies have reported the discovery of robust SC phases in twisted WSe2 (tWSe2) homobilayers [24, 25]. The phenomenology of these SC phases differ significantly from that would arise in doped Mott insulators, which call for further theoretical investigations, particularly regarding the possible pairing mechanism and topological properties of SC, as well as the nature of adjacent correlated insulators.
@@
Previous theoretical studies [9, 26, 27, 28, 16, 17, 19, 20, 21, 18, 22, 23] of tWSe2 homobilayers are largely based on the tight-binding description of the moiré Hubbard model with spin-dependent hopping phase tuned by the displacement field. This approach provides valuable real-space insights essential for understanding the correlated physics. However, such a simple one-band model cannot capture the potential nontrivial band topology [6, 29] and only applies to limited twist angles [9]. Thus, a direct Wannierization [29, 30, 31, 32, 33, 34] of the standard continuum model [6] with a proper set of low-energy bands appears to be a more suitable approach to start with.
@@
In this paper, we focus mainly on the experimental results in Ref. [24], where the SC phase is observed in a $3.65^{\circ}$ tWSe2 device at integer filling factor $\nu=-1$ under small displacement field, along with a correlated insulator phase at larger displacement field. Using the continuum model parameters provided in Ref. [29], we first construct a three-band tight-binding model for $3.65^{\circ}$ tWSe2 through direct Wannierization [29, 30, 31, 32, 33, 34]. In addition to the onsite Hubbard repulsion, our model includes the next-nearest-neighbor (NNN) attraction, which can arise effectively through electron-phonon coupling [35, 36, 37, 38] or fluctuations of the neighboring insulating phase [39, 40, 41]. We then perform a thorough mean-field analysis to the interacting model and obtain the phase diagram at filling $\nu=-1$ under realistic interaction strengths, concluding that the SC phase is consistent with inter-valley pairing with mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave symmetry, featuring non-trivial topology under Altland-Zirnbauer A-class with Chern number $C=\pm 2$, and the correlated insulator phase has $120^{\circ}$ antiferromagnetic (AFM) order. Additionally, the mixed symmetry character of the SC phase near zero displacement field predicted by our model has experiment observable distinctions from that of moiré Hubbard model [42, 43, 16, 17, 18] due to the absence of emergent spin-valley SU(2) symmetry [26, 27], providing new insights into the nature of the moiré TMD systems.
@@
Continuum model and symmetry analysis.—We begin with the standard low-energy continuum model description of tWSe2 [6]. The non-interacting continuum Hamiltonian for $K$ valley is given by:
@@
$\Delta_{\pm}(\bm{r})$ =±$=\pm\frac{V_{z}}{2}+2V\sum_{j=1,3,5}\cos\left(\bm{g}_{j}\cdot\bm{r}\pm\psi\right)$ (2)
@@
and the Hamiltonian for $-K$ valley is related by time-reversal symmetry $\mathcal{T}$. Throughout this paper, we adopt the experimental relevant twist angle $\theta=3.65^{\circ}$ [24], continuum model parameters derived by large scale ab initio simulations ($V$, $\psi$, $w$) = ($9$ meV, $128^{\circ}$, $18$ meV) [29] and the effective mass $m^{*}=0.45m_{e}$ [44]. The moiré lattice constant is given by $a_{M}\approx a_{0}/\theta$ with $a_{0}=3.317$Å [45].
In the absence of displacement field $V_{z}=0$, tWSe2 system has $D_{3}$ point group symmetry generated by threefold rotation $C_{3z}$ and twofold rotation $C_{2y}$. Additionally, due to the lowest order harmonic approximation we adopted, the continuum model has additional pseudo-inversion symmetry $\mathcal{I}$ with $\sigma_{x}H_{K}(\bm{r})\sigma_{x}=H_{K}(-\bm{r})$, enlarging the point group symmetry to $D_{3d}$. When a finite displacement field $V_{z}\neq 0$ is applied, the point group symmetry is reduced from $D_{3d}$ to $C_{3v}$, breaking all symmetries that interchange the two layers. Apart from point group symmetries, the system also exhibits U(1) spin-valley symmetry and time reversal symmetry $\mathcal{T}$ for both $\mathcal{V}_{z}=0$ and $\mathcal{V}_{z}\neq 0$. These symmetry considerations are crucial for the subsequent construction of tight binding model and the classification of SC pairing symmetries.
Wannier functions and tight-binding model.—To study the low-energy physics, we first focus on the top moiré valence bands of the tWSe2 continuum model at $V_{z}=0$. The Chern numbers for each valley share the same sign up to top five moiré bands, precluding any low-energy real-space description with correct band topology due to Wannier obstructions. Therefore, we focus on accurately reproducing the Chern numbers of the top two moiré bands, and construct a minimal three-band tight-binding model with Chern numbers ($1,1,-2$) for $K$ valley. Following Refs. [29, 30, 31], we construct sufficiently localized Wannier functions using their layer polarization properties and the location of Wannier centers. The resulting Wannier functions are shown in Fig. 1(b), where the A and B orbitals are centered at XM and MX regions (±$(\pm a_{M}/\sqrt{3},0)$ with opposite layer polarization, and the (more extended) C orbital is centered at MM region $(0,0)$ with layer hybridization. Detailed symmetry analysis [31] reveals that the (A, B, C) orbitals for the $K$ valley have $C_{3z}$ eigenvalues ($e^{-i2\pi/3}$, $e^{-i2\pi/3}$, $1$), and $C_{2y}\mathcal{T}$ or $\mathcal{I}$ symmetries interchange A and B orbitals but leave C orbital invariant. Using these Wannier functions, we can construct the following tight-binding model:
@@
where $i,j$ are unit cell indexes, $\alpha,\beta$ are sublattice indexes (A, B, C), and $\sigma$ is spin-valley index $\uparrow$ (or $K$), $\downarrow$ (or $-K$). By keeping up to fifth-nearest-neighbor hopping parameters [46], Fig. 1(c) illustrates the close matching between the band structure of the tight-binding model and the continuum model, especially for the top moiré valence band.
Figure 1: (a) Mini Brillouin zone formed by a small twist angle $\theta$ between two layers. (b) Density distribution of Wannier functions on two layers, with unit cell shown as dashed line. (c) Comparison of the band structure between continuum model (dashed line) and tight binding model (solid line), where colors representing sublattice components.
@@
where $\mathcal{V}_{z}$ is the energy difference between A and B sublattices induced by the displacement field, and $n_{i\alpha}=\sum_{\sigma}c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$ is the density operator. Since the energy expectation value of displacement field on the C sublattice can be chosen as 00, we neglect terms involving it in $H_{D}$ [47]. The Fermi surfaces of filling factor $\nu=-1$ with $\mathcal{V}_{z}=0$, $5$, $15$ and $25$ meV are shown in Fig. 2(a), where the Fermi surface of spin up and down are split by the displacement field $\mathcal{V}_{z}$, but related by the time reversal symmetry $\mathcal{T}$. Fig. 2(b) also illustrates the Fermi surface density of states (DOS) as a function of $\mathcal{V}_{z}$.
Figure 2: (a) Fermi surfaces of free Hamiltonian at different displacement fields, with colors indicating the sublattice components, and thickness representing the DOS. The spin of the Fermi surfaces are labeled as solid arrows. When $\mathcal{V}_{z}=0$ meV, where spin up and down Fermi surfaces coincide, only spin down Fermi surface is shown. The approximate nesting wave vector $\bm{Q}$ between spin up and down Fermi surfaces are illustrated as dashed arrow. (b) The Fermi surface DOS of A, B, C sublattices as a function of displacement field $\mathcal{V}_{z}$.
@@
where $\alpha=A,B,C$ labels the sublattice, and $n_{i\alpha\sigma}=c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$. We take $U_{A}=U_{B}>U_{C}$ due to the Wannier orbital distributions as discussed previously. The positive Hubbard interaction can typically lead to magnetic ordering, providing a promising explanation of the correlated insulator phase.
Due to the time reversal symmetry $\mathcal{T}$ of the systems, the Fermi surface features Cooper instability, namely SC instabilities can occur even with infinitesimal attractions, which can be effectively induced by electron-boson (phonon, magnon, etc.) couplings. Considering that the Fermi surface shown in Fig. 2(a) favor pairing within the same sublattices [46] and has almost no C-orbital component, a natural choice is to consider the attractions on NNN sites of the same A or B sublattices:
$H_{V}=-V\sum_{i}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in\text{NNN}}n_{i+\delta\alpha}n_{i\alpha}$ (7)
where $V$ is the strength of NNN attraction, $\alpha$ is sublattice index of A and B, and $\delta$ represents one of the three NNN bond directions $120^{\circ}$ apart. Besides, we have also considered the effects of nearest-neighbor (NN) attraction between A and B sublattices in the Supplemental Material [46], where we have shown that it is less dominant. The interacting model which we should consider to describe the main physics of tWSe2 is denoted as $H=H_{0}+H_{D}+H_{U}+H_{V}.$
Figure 3: (a) The NNN SC order parameters $\Delta_{AA}$ and $\Delta_{BB}$, where the A, B, C sites are labeled as red, green, blue dots. (b) The pairing form factor of $s$-wave, $f$-wave, $p+ip$-wave, and $d-id$-wave on NNN bonds, with the irreducible representation labeled. (c) $120^{\circ}$ AFM pattern on a certain type of sublattice with wave vector ±$\pm Q$. (d) The folded Brillouin zone induced by the AFM order.
Mean-field analysis of SC and AFM— We start with the mean-field analysis of SC instabilities by decoupling the NNN attraction $H_{V}$ in the SC channel as:
$H_{V}\approx-V\sum_{i\sigma\sigma^{\prime}}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\bigg{(}\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}^{*}(\delta)c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}$
@@
where $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)=\langle c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}\rangle$ is the spatially uniform real-space pairing order parameter on NNN bond $\delta$, as shown in Fig. 3(a). To further determine the ansatz of the pairing order parameter $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)$, we analyze the possible pairing symmetries, which should be classified by the irreducible representations of the symmetry group. Since the addition of displacement field $\mathcal{V}_{z}$ preserves time reversal symmetry $\mathcal{T}$ but breaks pseudo-inversion symmetry $\mathcal{I}$, we only consider $S_{z}=0$ sector for the U(1) spin-valley symmetry (i.e. inter-valley pairing). And these $S_{z}=0$ pairings can be further categorized into singlet pairing $\Delta_{\alpha}^{S}(\delta)=(\tilde{\Delta}_{\alpha\uparrow\downarrow}(\delta)-\tilde{\Delta}_{\alpha\downarrow\uparrow}(\delta))/\sqrt{2}$ and triplet pairing $\Delta_{\alpha}^{T}(\delta)=(\tilde{\Delta}_{\alpha\uparrow\downarrow}(\delta)+\tilde{\Delta}_{\alpha\downarrow\uparrow}(\delta))/\sqrt{2}$ [48]. For the point group symmetry $C_{3v}$ under finite displacement field, a classification of pairing symmetries based on momentum space continuum model has been provided in Ref. [49]. Switching to our tight-binding description, it is straightforward to show that only $A_{1}$ (mixing $s$ and $f$-wave) and $E$ (mixing $(p_{x},p_{y})$ and $(d_{xy},d_{x^{2}-y^{2}})$-wave) representations are possible for NNN pairing. The SC pairing form factors, focusing only on chiral SC in $E$ representation, are illustrated in Fig. 3(b), and we leave the discussion of nematic SC in $E$ representation in the Supplemental Material [46], where we have shown it is subdominant.
@@
$H_{U}=-\sum_{i\alpha}U_{\alpha}\left(m_{i\alpha}c_{i\alpha\downarrow}^{\dagger}c_{i\alpha\uparrow}+m_{i\alpha}^{*}c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\right)$ (9)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{i\alpha}U_{\alpha}|m_{i\alpha}|^{2}+\frac{1}{2}\sum_{i\alpha\sigma}U_{\alpha}n_{i\alpha\sigma}$
@@
where $m_{i\alpha}=\langle c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\rangle=\langle S_{i\alpha}^{x}\rangle+i\langle S_{i\alpha}^{y}\rangle$ is the complex order parameter representing the in-plane magnetization. We further constraint the form of $m_{i\alpha}$ by noting that, as shown in Fig. 2(a), although the Fermi surface deforms as the displacement field $\mathcal{V}_{z}$ changes, an approximate nesting between spin up and spin down Fermi surfaces persist. Moreover, the nesting wave vectors are relatively close to the commensurate ones ±$\pm\bm{Q}=(0,\pm 4\pi/3a_{M})$ over a wide range of displacement field $\mathcal{V}_{z}$, giving raise to the $120^{\circ}$ AFM order $m_{i\alpha}=m_{\alpha}^{+}e^{iQ\cdot R_{i}}+m_{\alpha}^{-}e^{-iQ\cdot R_{i}}$ with two possible chiralities [50] as shown in Fig. 3(c). Such AFM order breaks translation symmetry of the system, folding the Brillouin zone as shown in Fig. 3(d), with the new filling factor in terms of the folded Brillouin zone becoming $\tilde{\nu}=-3$. Depending on details of the system, the ground state can be either metallic or insulating [51] due to the nesting of ±$\pm\bm{Q}$ is non-perfect.
@@
We then perform mean-field calculations for SC and AFM independently, using intermediate Hubbard repulsion $U_{A}=U_{B}=35$ meV, $U_{C}=25$ meV and two different NNN attractions $V=10$ meV or $V=12.5$ meV as representing parameters. The resulting phase diagrams for tuning displacement field $\mathcal{V}_{z}$ at filling $\nu=-1$ are summarized in Fig. 4(e) and (f), respectively. And we leave more detailed mean-field derivations and discussions in the Supplemental Material [46].
Figure 4: (a) The energy gain per hole of the ordered phases compared to the symmetric phase. (b) SC orders for $V_{2}=10$ meV on one specific NNN bond $\delta_{0}$ under the gauge described in the main text. The sign of order parameters correspond to the sign of real(imaginary) part of singlet(triplet) pairings. (c) The AFM mean-field charge gap at filling factor $\tilde{\nu}=-3$ in the folded Brillouin zone. (d) Magnitude of the AFM orders on A, B, C sublattices. (e, f) The phase diagrams for $V_{2}=10$ meV and $V_{2}=12.5$ meV respectively, indicating that under mean-field framework, the SC to AFM-I transition can either have an intermediate AFM-M phase or occur directly.

### v2

@@
thanks: These two authors contribute equally in this workthanks: These two authors contribute equally in this work
Theory of Topological Superconductivity and Antiferromagnetic Correlated Insulators in Twisted Bilayer WSe2
Chuyi Tuo Institute for Advanced Study, Tsinghua University, Beijing 100084, China    Ming-Rui Li Institute for Advanced Study, Tsinghua University, Beijing 100084, China Department of Physics, Princeton University, Princeton, New Jersey 08544, USA    Zhengzhi Wu Institute for Advanced Study, Tsinghua University, Beijing 100084, China Rudolf Peierls Centre for Theoretical Physics, Parks Road, Oxford, OX1 3PU, UK    Wen Sun Institute for Advanced Study, Tsinghua University, Beijing 100084, China    Hong Yao yaohong@tsinghua.edu.cn Institute for Advanced Study, Tsinghua University, Beijing 100084, China
(October 31, 2025)
I Abstract
@@
Since the very recent discovery of unconventional superconductivity in twisted WSe2 homobilayers at filling $\nu=-1$, considerable interest has arisen in revealing its mechanism. In this paper, we developed a three-band tight-binding model with non-trivial band topology by direct Wannierization of the low-energy continuum model. Incorporating both onsite Hubbard repulsion and next-nearest-neighbor attraction, we then performed a mean-field analysis of the microscopic model and obtained a phase diagram qualitatively consistent with the experiment results. For zero or weak displacement field, the ground state is a Chern number $C=\pm 2$ topological superconductor in the Altland-Zirnbauer A-class (breaking time-reversal but preserving total $S_{z}$ symmetry) with inter-valley pairing dominant in $d_{xy}\pm id_{x^{2}-y^{2}}$-wave (mixing with a subdominant $p_{x}\mp ip_{y}$-wave) component. For a relatively strong displacement field, the ground state is a correlated insulator with the $120^{\circ}$ antiferromagnetic order. Our results provide new insights into the nature of the twisted WSe2 systems and suggest the need for further theoretical and experimental explorations.
II Introduction
@@
In recent years, twisted van der Waals moiré superlattices [1, 2, 3, 4, 5] have garnered significant interest following the groundbreaking discovery of unconventional superconductivity (SC) [6, 7, 8, 9] in magic-angle twisted bilayer graphene [10], which inspired numerous theoretical investigations on its microscopic origin [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Beyond graphene-based systems [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], twisted transition metal dichalcogenides (TMDs) [45, 46] are regarded as a promising alternative platform to investigate many-body physics. Due to their high tunability, twisted TMDs can host a wide variety of exotic phases, including correlated insulators [47, 48, 49, 50, 51, 52, 53], integer and fractional quantum anomalous Hall states [54, 55, 56, 57, 58], and also integer and fractional quantum spin Hall states [59]. However, the experimental realization of SC in twisted TMD systems has remained elusive [49], though many theoretical works have suggested that SC should develop in such systems [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], typically through doping the correlated insulators. Very recently, two independent studies have reported the discovery of robust SC phases in twisted WSe2 (tWSe2) homobilayers [71, 72]. The phenomenology of these SC phases differ significantly from that would arise in doped Mott insulators, which calls for further theoretical investigations, particularly regarding the possible pairing mechanism and topological properties of SC, as well as the nature of adjacent correlated insulators.
@@
Previous theoretical studies [49, 73, 74, 75, 60, 61, 67, 63, 64, 62, 65, 66, 76] of tWSe2 homobilayers are largely based on the tight-binding description of the single-band moiré Hubbard model with spin-dependent hopping phase tuned by the displacement field. By formulating the problem in real space, this approach captures the locality of screened interactions and facilitates a clearer real-space insights essential for understanding the correlated physics. However, such a simple single-band model cannot capture the potential nontrivial band topology [46, 77, 78, 79, 80] and only applies to limited twist angles [49]. Thus, a direct Wannierization [80, 81, 82, 83, 84, 85, 86] of the standard continuum model [46] with a proper set of low-energy bands appears to be a more suitable approach to start with.
@@
In this paper, we focus mainly on the experimental results in Ref. [71], where the SC phase is observed in a $3.65^{\circ}$ tWSe2 device at integer filling factor $\nu=-1$ under a small displacement field, along with a correlated insulator phase in a larger displacement field. Using the continuum model parameters provided in Ref. [80], we first construct a three-band tight-binding model for $3.65^{\circ}$ tWSe2 through direct Wannierization [80, 81, 82, 83, 84, 85, 86]. In addition to the onsite Hubbard repulsion, our model includes the next-nearest-neighbor (NNN) attraction, which can arise effectively through electron-phonon coupling [87, 88, 11, 12, 13, 14, 15, 16, 17, 21, 22, 23, 24, 25, 18, 19, 20] or purely electronic mechanisms [89, 90, 91, 92, 93, 94, 95, 96]. By keeping these key interaction terms, our model can capture the essential qualitative physics of the experiment [71] and offer valuable real-space insights into the twisted WSe2 system. We then perform a thorough mean-field analysis to the interacting model and obtain the phase diagram at filling $\nu=-1$ under realistic interaction strengths, concluding that the SC phase is consistent with inter-valley pairing with mixed $d_{xy}\pm id_{x^{2}-y^{2}}$ and $p_{x}\mp ip_{y}$-wave symmetry, featuring nontrivial topology under Altland-Zirnbauer A-class [97, 98, 99] with Chern number $C=\pm 2$, and the correlated insulator phase has $120^{\circ}$ antiferromagnetic (AFM) order. Additionally, the mixed symmetry character of the SC phase near zero displacement field predicted by our model has experiment observable distinctions from that of moiré Hubbard model [49, 73, 74, 75, 60, 61, 67, 63, 64, 62, 65, 66, 76, 100, 101] due to the absence of emergent spin-valley SU(2) symmetry [73, 74], providing new insights into the nature of the moiré TMD systems.
III Results
@@
Continuum model and symmetry analysis.—We begin with the standard low-energy continuum model description of tWSe2 [46]. The non-interacting continuum Hamiltonian for the $K$ valley is given by:
@@
$\Delta_{\pm}(\bm{r})$ $=\pm\frac{V_{z}}{2}+2V\sum_{j=1,3,5}\cos\left(\bm{g}_{j}\cdot\bm{r}\pm\psi\right)$ (2)
@@
and the Hamiltonian for K-K valley is related by time-reversal symmetry $\mathcal{T}$. Throughout this paper, we adopt the experimental relevant twist angle $\theta=3.65^{\circ}$ [71], continuum model parameters derived by large scale ab initio simulations ($V$, $\psi$, $w$) = ($9$ meV, $128^{\circ}$, $18$ meV) [80] and the effective mass $m^{*}=0.45m_{e}$ [102]. The moiré lattice constant is given by $a_{M}\approx a_{0}/\theta$ with $a_{0}=3.317$Å [103].
In the absence of displacement field $V_{z}=0$, tWSe2 system has $D_{3}$ point group symmetry generated by threefold rotation $C_{3z}$ and twofold rotation $C_{2y}$. Additionally, due to the lowest order harmonic approximation we adopted, the continuum model has additional pseudo-inversion symmetry $\mathcal{I}$ with $\sigma_{x}H_{K}(\bm{r})\sigma_{x}=H_{K}(-\bm{r})$, enlarging the point group symmetry to $D_{3d}$. When a finite displacement field $V_{z}\neq 0$ is applied, the point group symmetry is reduced from $D_{3d}$ to $C_{3v}$, breaking all symmetries that interchange the two layers. Apart from point group symmetries, the system also exhibits U(1) spin-valley symmetry and time-reversal symmetry $\mathcal{T}$ for both $V_{z}=0$ and $V_{z}\neq 0$. These symmetry considerations are crucial for the subsequent construction of tight binding model and the classification of SC pairing symmetries.
Wannier functions and tight-binding model.—To study the low-energy physics, we first focus on the top moiré valence bands of the tWSe2 continuum model at $V_{z}=0$. The Chern numbers for each valley share the same sign up to top five moiré bands, precluding any low-energy real-space description with correct band topology due to Wannier obstructions. Therefore, we focus on accurately reproducing the Chern numbers of the top two moiré bands, and construct a minimal three-band tight-binding model with Chern numbers (1,1,21,1,-2) for $K$ valley. Following Refs. [80, 81, 82, 83], we construct sufficiently localized Wannier functions using their layer polarization properties and the location of Wannier centers. The resulting Wannier functions are shown in Fig. 1(b), where the A and B orbitals are centered at XM and MX regions $(\pm a_{M}/\sqrt{3},0)$ with opposite layer polarization, and the C orbital is centered at MM region $(0,0)$ with layer hybridization. Detailed symmetry analysis [82] reveals that the (A, B, C) orbitals for the $K$ valley have $C_{3z}$ eigenvalues ($e^{-i2\pi/3}$, $e^{-i2\pi/3}$, $1$), and $C_{2y}\mathcal{T}$ or $\mathcal{I}$ symmetries interchange A and B orbitals but leave C orbital invariant. Using these Wannier functions, we can construct the following tight-binding model:
@@
where $i,j$ are unit cell indexes, $\alpha,\beta$ are sublattice indexes (A, B, C), and $\sigma$ is spin-valley index $\uparrow$ (or $K$), $\downarrow$ (or K-K). By keeping up to fifth-nearest-neighbor hopping parameters (see Sec. I of the Supplementary Information [104]), Fig. 1(c) illustrates the close matching between the band structure of the tight-binding model and the continuum model, especially for the top moiré valence band.
@@
where $\mathcal{V}_{z}$ is the energy difference between A and B sublattices induced by the displacement field, and $n_{i\alpha}=\sum_{\sigma}c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$ is the density operator. Since the energy expectation value of displacement field on the C sublattice can be chosen as 0, we neglect terms involving it in $H_{D}$ (see Sec. II of the Supplementary Information [104] for a more detailed discussion of the displacement field effects). The Fermi surfaces of filling factor $\nu=-1$ with $\mathcal{V}_{z}=0$, $5$, $15$ and $25$ meV are shown in Fig. 2(a), where the Fermi surface of spin up and down are split by the displacement field $\mathcal{V}_{z}$, but related by the time-reversal symmetry $\mathcal{T}$. Fig. 2(b) also illustrates the Fermi surface density of states (DOS) as a function of $\mathcal{V}_{z}$.
@@
where $\alpha=A,B,C$ labels the sublattice, and $n_{i\alpha\sigma}=c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$. The positive Hubbard interaction can typically lead to magnetic ordering, providing a promising explanation of the correlated insulator phase. Longer-range Coulomb repulsions are neglected mainly because their magnitudes are much smaller (see Sec. IV of the Supplementary Information [104]).
Due to the time-reversal symmetry $\mathcal{T}$ of the systems, the Fermi surface features Cooper instability; namely, SC instabilities can occur even with infinitesimal attractions. Considering that the Fermi surface shown in Fig. 2(a) favor pairing within the same sublattices and has almost no C-orbital component, a natural choice is to consider the attractions on NNN sites of the same A or B sublattices:
$H_{V_{2}}=-V_{2}\sum_{i}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in\text{NNN}}n_{i+\delta\alpha}n_{i\alpha}$ (7)
where $V_{2}$ is the strength of NNN attraction, $\alpha$ is sublattice index of A and B, and $\delta$ represents one of the three NNN bond directions $120^{\circ}$ apart. Besides, we have also considered the effects of nearest-neighbor (NN) attraction between A and B sublattices in Sec. VI of the Supplementary Information [104], where we have shown that it is less dominant. The relatively local NNN (or NN) attraction can be understood by a random phase approximation (RPA) analysis (see Sec. V of the Supplementary Information [104]). Therefore, the interacting model which we should consider to describe the main physics of tWSe2 is denoted as $H=H_{0}+H_{D}+H_{U}+H_{V_{2}}$.
Mean-field analysis of SC and AFM— We start with the mean-field analysis of SC instabilities by decoupling the NNN attraction $H_{V_{2}}$ in the SC channel as:
$H_{V_{2}}\approx-V_{2}\sum_{i\sigma\sigma^{\prime}}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\bigg(\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}^{*}(\delta)c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}$
@@
where $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)=\langle c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}\rangle$ is the spatially uniform real-space pairing order parameter on NNN bond $\delta$, as shown in Fig. 3(a). To further determine the ansatz of the pairing order parameter $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)$, we analyze possible pairing symmetries, which should be classified by the irreducible representations of the symmetry group. Since the addition of displacement field $\mathcal{V}_{z}$ preserves time-reversal symmetry $\mathcal{T}$ but breaks pseudo-inversion symmetry $\mathcal{I}$, we only consider the $S_{z}=0$ sector for the U(1) spin-valley symmetry (i.e. inter-valley pairing). And these $S_{z}=0$ pairings can be further categorized into singlet pairing $\Delta_{\alpha}^{S}(\delta)=(\tilde{\Delta}_{\alpha\uparrow\downarrow}(\delta)-\tilde{\Delta}_{\alpha\downarrow\uparrow}(\delta))/\sqrt{2}$ and triplet pairing $\Delta_{\alpha}^{T}(\delta)=(\tilde{\Delta}_{\alpha\uparrow\downarrow}(\delta)+\tilde{\Delta}_{\alpha\downarrow\uparrow}(\delta))/\sqrt{2}$. For the point group symmetry $C_{3v}$ under finite displacement field, it is straightforward to show that only $A_{1}$ (mixing $s$ and $f$-wave) and $E$ (mixing $(p_{x},p_{y})$ and $(d_{xy},d_{x^{2}-y^{2}})$-wave) representations are possible for NNN pairing. The SC pairing form factors, focusing only on chiral SC in the $E$ representation, are illustrated in Fig. 3(b), and we leave the discussion of nematic SC in the $E$ representation in Sec. VI of the Supplementary Information [104], where we have shown it is subdominant.
@@
$H_{U}=-\sum_{i\alpha}U_{\alpha}\left(m_{i\alpha}c_{i\alpha\downarrow}^{\dagger}c_{i\alpha\uparrow}+m_{i\alpha}^{*}c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\right)$
$~~~~~~~~~+\sum_{i\alpha}U_{\alpha}|m_{i\alpha}|^{2}+\frac{1}{2}\sum_{i\alpha\sigma}U_{\alpha}n_{i\alpha\sigma}$ (9)
@@
where $m_{i\alpha}=\langle c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\rangle=\langle S_{i\alpha}^{x}\rangle+i\langle S_{i\alpha}^{y}\rangle$ is the complex order parameter representing the in-plane magnetization. We further constraint the form of $m_{i\alpha}$ by noting that, as shown in Fig. 2(a), although the Fermi surface deforms as the displacement field $\mathcal{V}_{z}$ changes, an approximate nesting between spin up and spin down Fermi surfaces persist. Moreover, the nesting wave vectors are relatively close to the commensurate ones $\pm\bm{Q}=(0,\pm 4\pi/3a_{M})$ over a wide range of displacement field $\mathcal{V}_{z}$, giving raise to the $120^{\circ}$ AFM order $m_{i\alpha}=m_{\alpha}^{+}e^{iQ\cdot R_{i}}+m_{\alpha}^{-}e^{-iQ\cdot R_{i}}$ with two possible chiralities as shown in Fig. 3(c), which are generally non-degenerate under finite $\mathcal{V}_{z}$. Such AFM order breaks translation symmetry of the system, folding the Brillouin zone as shown in Fig. 3(d), with the new filling factor in terms of the folded Brillouin zone becoming $\tilde{\nu}=-3$. Depending on details of the system, the ground state can be either metallic or insulating due to the nesting of $\pm\bm{Q}$ is non-perfect.
@@
We then perform mean-field calculations for SC and AFM at filling $\nu=-1$ independently, using Hubbard repulsion $U_{A}=U_{B}=U_{C}=37.5$ meV and two different NNN attractions $V_{2}=10$ meV or $V_{2}=12.5$ meV as representing parameters. These interacting parameters can be estimated either by comparing the mean-field results with experimental observations, or by directly expanding the gate-screened Coulomb interaction onto Wannier functions (see Sec. III and Sec. IV of the Supplementary Information [104]). The resulting phase diagrams for tuning the displacement field $\mathcal{V}_{z}$ at filling $\nu=-1$ are summarized in Figs. 4(e) and (f), respectively. We leave more detailed mean-field derivations and discussions in Sec. VI and Sec. VII of the Supplementary Information [104].